Suppose you have 5 distinguishable dice. Calculate the size of the sample space.
List all the outcomes which satisfy the event "all five dice are the same".
Calculate the probability of thus event occurring. ( in other worlds calculate the probability of getting a yahtzee on one roll) leave answer as a fraction reduced to simplest terms

Answer by user_dude2008(1862)

size of sample space: 6^5 = 7776


List all the outcomes which satisfy the event "all five dice are the same".
{1,1,1,1,1}, {2,2,2,2,2}, ..., {5,5,5,5,5}, {6,6,6,6,6} ---> 6 ways


Calculate the probability of thus event occurring.

6/7776 = 1/1296

Answer: 1/1296